An engine raises 100kg of water through a height of 60m in 20s. what is the power of the engine​

1
Expert's answer
2022-01-06T10:46:41-0500
"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{100\ kg\times9.8\ \dfrac{m}{s^2}\times60\ m}{20\ s}=2940\ W."